OpenStudy (anonymous):

I have the answer can someone please check me...... find the derivative of c(t)=3te^5t

OpenStudy (anonymous):

c(t) = 3t* (e^t)^5 use product rule

OpenStudy (anonymous):

I got c'(t)=15e^5t+3te^5t is that correct

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

What did I do wrong

OpenStudy (anonymous):

c'(t) = e^(5t) *3 + 5*e^(4t)*e^t = 3e^(5t) + 5e^(5t) = 8e^(5t)

OpenStudy (anonymous):

I don't have that answer in any of my multiple choice answers

OpenStudy (anonymous):

\[c(t) = 3t*e ^{5t}\] is this ur question?

OpenStudy (anonymous):

no\[c(t)=3te ^{5t}\]

OpenStudy (anonymous):

Is the answer \[\frac{d}{dt}c(t)=3e^{5t}+15te^5t\]?

OpenStudy (anonymous):

Then @agostino is correct

OpenStudy (anonymous):

That's what I got. @Yahoo! You don't reduce the power of the exponential of e.

OpenStudy (anonymous):

yes I have that for one of my answers

OpenStudy (anonymous):

i found my mistake

OpenStudy (anonymous):

c'(t) = 3e^(5t) + 5e^(5t)*3t = 3e^(5t) + 15te^(5t)

OpenStudy (anonymous):

ok thanks everyone